OpenStudy (anonymous):

What is the distance between points (6, 5) and (–12, 17)? Round to the nearest tenth. 8.9 14.7 18.2 21.6

OpenStudy (anonymous):

@Nali

OpenStudy (nali):

so you got the ohter question?

OpenStudy (anonymous):

yeah he gave me the answer

OpenStudy (anonymous):

\[d = \sqrt{(x_{1} - x_{2})^2 + (y_1 - y_2)^2}\]

OpenStudy (nali):

for this one use the distance formula http://math.about.com/library/bldistance.htm

OpenStudy (nali):

yeah and just plug in the x and y values like before

OpenStudy (anonymous):

\[d = \sqrt{(6) - (-12))^2 + ((5) - (17))^2}\]\[d = \sqrt{18^2 + 12^2}\]\[d = ?

OpenStudy (anonymous):

ok so is it 21.6

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thank you ;)

OpenStudy (anonymous):

np :)